OpenStudy (anonymous):

I don't understand how to do this assignment that is posted below please help me!!

OpenStudy (anonymous):

assignment

OpenStudy (anonymous):

"A compact car will drive three miles east, five miles west, and then 13 miles east again. The car" Can you tell me how many miles she/he will drive?

OpenStudy (anonymous):

@Lissa1199

OpenStudy (anonymous):

Okay well. A compact car will drive three miles east, five miles west, and then 13 miles east again. The car so you add. 3 + 5 + 13 = x Write down your answer to 'x' on a piece of paper. There is 1/4 tank of gas so divide 40/4 = n Now, n - x = ? If the answer is negative, you don't have enough gasoline. If the answer is positive, you do have enough gasoline.

OpenStudy (anonymous):

Any questions?

OpenStudy (anonymous):

no thank you

OpenStudy (anonymous):

@SmokeyRiver could you help me on number 2. I have the first part 10+15+5=30

OpenStudy (anonymous):

Sure, what's the question

OpenStudy (anonymous):

Scenario 2: A van will drive ten miles north, 15 miles south, and then five miles north again. The van gets 33 miles per gallon, and there is one gallon of gas left in the tank. Compute the relative distance and the total distance. Then decide if there is enough fuel for the trip.

OpenStudy (anonymous):

Relative distance is 0. total distance is 30 miles. It's a 30 mile trip and you can travel 33 miles left with the tank of gas. So yes, there is enough fuel for the trip.

OpenStudy (anonymous):

A sedan will drive ten miles north, 15 miles east, 13 miles south, 15 miles west, and 22 miles north. The sedan gets 20 miles per gallon, and there are five gallons of gas left in the tank. Compute the relative distance and the total distance. Then decide if there is enough fuel for the trip. can this be solved the same way

OpenStudy (anonymous):

Sometimes math makes things sound more complicated then they are. Total distance is the total distance, so yea, 30 miles. Relative distance is how far they got. So instead of north/south I'll say forward/backword Instead of miled I'll say feet. If you walk 10 feet forward Then you walk back 15 feet Then you walk forward again, 5 feet. You would have gotten no where.

OpenStudy (anonymous):

So the relative distance would be 0

OpenStudy (anonymous):

" A sedan will drive ten miles north, 15 miles east, 13 miles south, 15 miles west, and 22 miles north. " A total of 75 miles. 20mpg on 5gallons would be 20*5 which equals?

OpenStudy (anonymous):

100

OpenStudy (anonymous):

100 - 75 = ? and is it positive or negative? Positive - You have enough gas Negative - You need more gas

OpenStudy (anonymous):

positive

OpenStudy (anonymous):

And I the relative distance is 9. If you have graph paper I find that it's easier to use that. I just got one off of google images and filled it in with paint:

OpenStudy (anonymous):

Total distance: 75 Relative distance: 9

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

No problem. :D